There are 9 students in a class: 5 boys and 4 girls.
If the teacher picks a group of 4 at random, what is the probability that everyone in the group is a boy?

Answer by Boreal(15235)

(5/9)(4/8)(3/7)(2/6)=120/3024=0.0397
(5C4)/(9C4)=0.0397